OpenStudy (swag):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was 3/4 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are dimensions of the brand name television? Answer 6 inches by 12 inches 9 inches by 18 inches 16 inches by 32 inches 18 inches by 36 inches

OpenStudy (hoblos):

dimensions of the brand name television = x inches by y inches 12 = (3/4)x -----> x= 4*12/3 = 16 24= (3/4)y ------> y=4*24/3 = 32

OpenStudy (hoblos):

so 16 inches by 32 inches